How much water must be mixed with 4 pints of a solution that is 90% developer to product a mixture that is 21% developer?
Options
13 1/7 pt
13 1/3 pt
16 20/21 pt
17 1/7 pt

Answer by stanbon(75887)

How much water must be mixed with 4 pints of a solution that is 90% developer to product a mixture that is 21% developer?
------
Equation:
active + active = active
0.90*4 + 0*x = 0.21(4+x)
-------
90*4 + 0 = 21*4 + 21x
21x = 69*4
x = 13.1429 pints (amt. of water to add)
x = 13 1/7 pts
--------
Cheers,
Stan H.
-----------


Options
13 1/7 pt
13 1/3 pt
16 20/21 pt
17 1/7 pt